# NeuralSurv: Deep Survival Analysis with Bayesian Uncertainty Quantification

### v1

@@ 4.2 Local Linearization of the Bayesian Neural Network
This transformation is particularly advantageous when placing a GP prior on $g(\cdot;\boldsymbol{\theta})$, as it induces conjugacy in the model. Conjugacy is crucial for variational inference because it enables efficient computation of the ELBO (16), which involves taking expectations over the distribution of $\boldsymbol{\theta}$. However, when $g(\cdot;\boldsymbol{\theta})$ is a BNN, these expectations generally lack closed-form solutions, making exact Bayesian updates intractable. As a result, we seek to approximate $g(\cdot;\boldsymbol{\theta})$ in a way that retains the expressive power of neural networks while preserving Gaussian conjugacy to enable tractable inference.
@@
$g(t,\mathbf{x};\boldsymbol{\theta})$ $\approx g^{\text{lin}}(t,\mathbf{x};\boldsymbol{\theta}):=g(t,\mathbf{x};\boldsymbol{\theta}^{\star})+\mathbf{J}_{\boldsymbol{\theta}^{\star}}(t,\mathbf{x})^{\top}(\boldsymbol{\theta}-\boldsymbol{\theta}^{\star}),$ (17)
where $[\mathbf{J}_{\boldsymbol{\theta}}(t,\mathbf{x})]_{j}=\frac{\partial g(t,\mathbf{x};\boldsymbol{\theta})}{\partial\theta_{j}}$ is the Jacobian of the BNN with respect to the parameters $\boldsymbol{\theta}$. Following Immer2021 , we select $\boldsymbol{\theta}^{\star}=\boldsymbol{\theta}_{\text{MAP}}$ as the maximum a posteriori (MAP) estimate, which is defined as:
@@
where $p(\boldsymbol{\theta},\phi\mid\mathcal{D},\mathbf{X})$ is the posterior density defined in (7). By centering the linearization at $\boldsymbol{\theta}_{\text{MAP}}$, we ensure maximal approximation accuracy precisely where Bayesian inference is most sensitive — in the high-probability region of the posterior that dominates both parameter uncertainty quantification and predictive distributions. The procedure used to obtain the MAP estimates of (18) is detailed in Appendix F. Under the assumption of a Gaussian prior on the BNN parameters (4), the local linearization induces the GP prior
@@
$\mu(t,\mathbf{x})$ $:=g(t,\mathbf{x};\boldsymbol{\theta}_{\text{MAP}})+\mathbf{J}_{\boldsymbol{\theta}_{\text{MAP}}}(t,\mathbf{x})^{\top}(\mathbb{E}_{\boldsymbol{\theta}\sim p_{\boldsymbol{\theta}}}[\boldsymbol{\theta}]-\boldsymbol{\theta}_{\text{MAP}})$
$\kappa((t,\mathbf{x}),(t^{\prime},\mathbf{x}^{\prime}))$ $:=\mathbf{J}_{\boldsymbol{\theta}_{\text{MAP}}}(t,\mathbf{x})\mathbf{J}_{\boldsymbol{\theta}_{\text{MAP}}}(t^{\prime},\mathbf{x}^{\prime})^{\top}.$
Incorporating this approximation into our variational framework allows us to exploit Gaussian conjugacy for fast, closed-form updates, while still preserving the flexibility of neural networks. Concretely, we take a Taylor expansion of the ELBO around $g^{\text{lin}}$ and, by truncating at the lowest order term, obtain the simple approximation
@@
Our approach is analogous to the method introduced in (variational_inference_nonconjugate, , Section 3.2), where the authors apply Delta Method Variational Inference by approximating the ELBO around a fixed point in parameter space. In contrast, we extend this idea by approximating the ELBO around a reference function $g^{\text{lin}}$, rather than a fixed point.
@@ 4.3 Coordinate Ascent Variational Inference
We adopt a Coordinate Ascent Variational Inference (CAVI) approach, allowing us to draw on standard results from variational inference (see, e.g., (bishop_pattern_recognition, , Chapter 10.1)). In this framework, the optimal variational distributions are derived by maximizing the linearized ELBO, $\mathcal{L}_{\text{ELBO}}(g^{\text{lin}})$, with each distribution depending on the current state of the others. The algorithm proceeds by cyclically updating each variational distribution while keeping the others fixed. This iterative process progressively refines the optimal variational distributions, ultimately leading to the best possible approximation of the posterior distribution. A complete derivation of each optimal variational distribution is provided in Appendix G while the complete CAVI algorithm is presented in Appendix H.
@@
where $(\tilde{\alpha}^{(k)},\tilde{\beta})$ and $(\tilde{\boldsymbol{\mu}}^{(k)},\tilde{\boldsymbol{\Sigma}}^{(k)})$ are given in Appendix G.3 and G.4, respectively. At the $k^{\text{th}}$ iteration, the optimal update for the auxiliary parameters $\boldsymbol{\omega}$ is given by
@@
where $\tilde{c}_{i}^{(k)}$ is given in Appendix G.1. Finally, at the $k^{\text{th}}$ iteration, the optimal variational law $\mathbb{Q}_{\boldsymbol{\Psi}}^{(k)}$ is the probability measure under which each $\Psi_{i}$ ($i=1,\ldots,N$) is a marked Poisson process on $[0,y_{i}]\times\mathbb{R}_{+}$ with intensity function $\lambda_{i}^{\mathbb{Q},(k)}$, as given in Appendix G.2.
It is important to emphasize that we did not impose a specific form on the variational distributions — for example, we did not assume $q_{\boldsymbol{\theta}}(\boldsymbol{\theta})$ to be Gaussian. Instead, we derived our results by variationally minimizing the KL divergence over the full space of distributions. This contrasts with methods that fix a parametric form and use the reparameterization trick with Monte Carlo gradient estimates.
Finally, in Appendix I, we demonstrate that, by exploiting the Woodbury matrix identity, our inference updates require only $\mathcal{O}(m)$ time complexity ($m$ is the number of weights in the neural network architecture). This linear scaling renders our Bayesian framework feasible for contemporary large-scale deep neural architectures, which are well suited to model high-dimensional data.
@@ 5 Experiments
Details on the experimental setup, including dataset descriptions, hyperparameter tuning for the benchmark methods and evaluation metrics definitions are provided in Appendix J. Moreover, the implementation details for NeuralSurv are provided in Appendix K.
To comprehensively evaluate NeuralSurv, we compare its performance against the same set of benchmark models evaluated in DySurv Mesinovic2024 . These include: MTLR Yu2011 , DeepHit Lee2018 , DeepSurv Katzman2018 , Logistic Hazard Gensheimer2019 , CoxTime Kvamme2019 , CoxCC Kvamme2019 , PMF Kvamme2021 , PCHazard Kvamme2021 , BCESurv Kvamme2023 , and DySurv Mesinovic2024 . A detailed overview of these methods is provided in Appendix L and Table A1. Except for DySurv, which employs an autoencoder framework, we adopt the same neural-network architecture across all benchmark models and NeuralSurv to parameterize the hazard function. For DySurv, we use the original autoencoder architecture specified in its implementation.
We assess discriminative performance using the Antolini’s concordance index (C-index) Antolini2005 , and evaluate model calibration with the inverse probability of censoring weighting (IPCW) integrated Brier score (IBS) Graf1999 . The C-index evaluates how well a model performs by measuring the concordance between the rankings of the predicted event times and the true event times. The C-index ranges from 0 to 1, where higher values indicate better discriminative performance; a value of 0.5 corresponds to random guessing. Similar to the mean squared error, the Brier score (BS) assesses the accuracy of an estimated survival function at some time $t$. The IPCW are observation-specific weights that account for censoring in survival data, ensuring that the BS remains unbiased. The IPCW IBS is the integral of the IPCW BS over the observational period. The C-index and the IPCW IBS metrics are computed using the TorchSurv package Monod2024 .
@@ 5.1 Synthetic Data Experiment
In this section, we present experiments conducted on synthetic data. The experimental setup was inspired by teh_survival and constitutes a broadly applicable evaluation benchmark. We simulated the training sets with increasing sizes $N=25,50,100$ and $150$ samples where the event time was drawn from two distributions: $p_{0}(T)=\log\text{Normal}(3,0.8^{2})$ and $p_{1}(T)=\log\text{Normal}(3.5,1^{2})$. Each observation included a covariate indicating whether the event time was sampled from $p_{0}$ or $p_{1}$, along with three additional noisy covariates generated from a standard normal distribution. The censoring times were drawn from an exponential distribution with a rate of $0.025$ yielding an average censoring rate of 54% across the four synthetic datasets. The test set was generated using the same data-generating process, fixed to $100$ observations, and held constant across all experiments.
Figure 1 presents the true survival function alongside the predicted functions from NeuralSurv and the two top-performing benchmark models, selected based on IPCW IBS. Each panel represents a different training set size. As the number of training samples increases, the predicted survival functions match more closely the true survival function. The results show that NeuralSurv consistently ranks as the best method according to IPCW IBS, and its predictive accuracy improves with larger sample sizes. Beyond its competitive performance, NeuralSurv also provides Bayesian credible intervals, offering uncertainty estimates for survival probabilities, an important feature typically absent in benchmark models. Notably, these credible intervals appropriately narrow as more data becomes available, demonstrating well-calibrated uncertainty quantification. Corresponding C-index and IPCW IBS scores for all methods are reported in Table A2.
@@ 5.2 Real Survival Data Experiments
Table 1 presents the results on the held-out test sets for three representative datasets, while results for the remaining datasets are shown in Table A3. Across all eight datasets, NeuralSurv achieves the best IPCW IBS score on seven, demonstrating superior calibration performance compared to benchmark methods. This improvement can be attributed to the Bayesian framework, which naturally incorporates model uncertainty and provides better regularization in data-scarce settings.
Additional an ablation study using a larger training set of 250 observations is presented in Table A4. NeuralSurv continues to outperform benchmark methods under this setting in terms of calibration performance demonstrating the robustness of the method to training size. Furthermore, we also include results from traditional survival models, such as the Cox Proportional Hazards model Cox1972 , the Weibull Accelerated Failure Time model Carroll2003 , the Random Survival Forest Ishwaran2008 , and the Survival Support Vector Machine Polsterl2015 in Table A5. These models, reported, often achieve strong performance in data-scarce regimes. However, they are not designed to leverage high-dimensional or complex feature representations, which limits their applicability in modern deep learning contexts. Our focus remains on evaluating deep survival methods that can scale with data complexity, but we include these classical baselines for reference and completeness.
@@ 6 Conclusion
We propose the first fully Bayesian framework for deep survival analysis with time-varying covariate–risk relationships. On both synthetic and real-world datasets, in data-scarce regimes, our method consistently achieves better calibration than state-of-the-art deep survival models and matches or surpasses their discriminative performance, while offering fully Bayesian uncertainty quantification.
Despite its strengths, NeuralSurv relies on three key simplifying assumptions. First, we assume a sigmoidal hazard function, a choice shared by prior work (e.g.,  teh_survival ; Kim2018 ), which may not capture all risk dynamics. Second, our mean-field variational inference treats parameters $\phi$ and $\boldsymbol{\theta}$ as independent, ignoring posterior correlations. Third, we linearize the network around the MAP estimate to enforce conjugacy. In real-world settings, however, the true posterior can be multimodal and strongly correlated, so this local, factorized approximation may overlook secondary modes or misestimate joint uncertainty.
Concerning the computational efficiency of our method, the coordinate-ascent updates scale linearly with network size but still require full-dataset passes each iteration. For very large cohorts, this becomes a bottleneck. Extending the algorithm to use stochastic or mini-batch updates would preserve conjugacy benefits while improving scalability.

### v2

@@ 4.2 Local Linearization of the Bayesian Neural Network
This transformation is particularly advantageous when placing a GP prior on $g(\cdot;\boldsymbol{\theta})$, as it induces conjugacy in the model. Conjugacy is crucial for variational inference because it enables efficient computation of the ELBO (17), which involves taking expectations over the distribution of $\boldsymbol{\theta}$. However, when $g(\cdot;\boldsymbol{\theta})$ is a BNN, these expectations generally lack closed-form solutions, making exact Bayesian updates intractable. As a result, we seek to approximate $g(\cdot;\boldsymbol{\theta})$ in a way that retains the expressive power of NNs while preserving Gaussian conjugacy to enable tractable inference.
@@
$g(t,\mathbf{x};\boldsymbol{\theta})$ $\approx g^{\text{lin}}(t,\mathbf{x};\boldsymbol{\theta})=g(t,\mathbf{x};\boldsymbol{\theta}^{\star})+\mathbf{J}_{\boldsymbol{\theta}^{\star}}(t,\mathbf{x})^{\top}(\boldsymbol{\theta}-\boldsymbol{\theta}^{\star}),$ (18)
where $\mathbf{J}_{\boldsymbol{\theta}}(t,\mathbf{x})]_{j}=\frac{\partial g(t,\mathbf{x};\boldsymbol{\theta})}{\partial\theta_{j}}$ is the Jacobian of the BNN with respect to the parameters $\boldsymbol{\theta}$. Following Immer2021, we select $\boldsymbol{\theta}^{\star}=\boldsymbol{\theta}_{\text{MAP}}$ as the maximum a posteriori (MAP) estimate, which is defined as:
@@
where $p(\boldsymbol{\theta},\phi\mid\mathcal{D},\mathbf{X})$ is the posterior density defined in (7). By centering the linearization at $\boldsymbol{\theta}_{\text{MAP}}$, we ensure maximal approximation accuracy precisely where Bayesian inference is most sensitive: in the high-probability region of the posterior that dominates both parameter uncertainty quantification and predictive distributions. The procedure used to obtain the MAP estimates of (19) is detailed in Appendix F. Under the assumption of a Gaussian prior on the BNN parameters (5), the local linearization induces the GP prior
@@
$\mu(t,\mathbf{x})$ $=g(t,\mathbf{x};\boldsymbol{\theta}_{\text{MAP}})+\mathbf{J}_{\boldsymbol{\theta}_{\text{MAP}}}(t,\mathbf{x})^{\top}(\mathbb{E}_{\boldsymbol{\theta}\sim p_{\boldsymbol{\theta}}}[\boldsymbol{\theta}]-\boldsymbol{\theta}_{\text{MAP}})$
$\kappa((t,\mathbf{x}),(t^{\prime},\mathbf{x}^{\prime}))$ $=\mathbf{J}_{\boldsymbol{\theta}_{\text{MAP}}}(t,\mathbf{x})\mathbf{J}_{\boldsymbol{\theta}_{\text{MAP}}}(t^{\prime},\mathbf{x}^{\prime})^{\top}.$
Incorporating this approximation into our variational framework allows us to exploit Gaussian conjugacy for fast, closed‐form updates, while still preserving the flexibility of NNs. Concretely, we take a Taylor expansion of the ELBO around $g^{\text{lin}}$ and, by truncating at the lowest order term, obtain the simple approximation
@@
Our approach is analogous to the method introduced in (variational_inference_nonconjugate, Section 3.2), where the authors apply Delta Method Variational Inference by approximating the ELBO around a fixed point in parameter space. In contrast, we extend this idea by approximating the ELBO around a reference function $g^{\text{lin}}$, rather than a fixed point.
@@ 4.3 Coordinate Ascent Variational Inference
We adopt a Coordinate Ascent Variational Inference (CAVI) approach, allowing us to draw on standard results from variational inference (see, e.g., (bishop_pattern_recognition, Chapter 10.1)). In this framework, the optimal variational distributions are derived by maximizing the linearized ELBO, $\mathcal{L}_{\text{ELBO}}(g^{\text{lin}})$, with each distribution depending on the current state of the others. The algorithm proceeds by cyclically updating each variational distribution while keeping the others fixed. This iterative process progressively refines the optimal variational distributions, ultimately leading to the best possible approximation of the posterior distribution. A complete derivation of each optimal variational distribution is provided in Appendix G while the complete CAVI algorithm is presented in Appendix H.
@@
[…]
@@
where $\tilde{c}_{i}^{(k)}$ is given in Appendix G.1. Finally, at the $k^{\text{th}}$ iteration, the optimal variational law $\mathbb{Q}_{\boldsymbol{\Psi}}^{(k)}$ is the probability measure under which each $\Psi_{i}$ ($i=1,\ldots,N$) is a marked Poisson process on $0,y_{i}]\times\mathbb{R}_{+}$ with intensity function $\lambda_{i}^{\mathbb{Q},(k)}$, as given in Appendix G.2.
It is important to emphasize that we did not impose a specific form on the variational distributions; for example, we did not assume $q_{\boldsymbol{\theta}}(\boldsymbol{\theta})$ to be Gaussian. Instead, we derived our results by minimizing the KL divergence over the full space of distributions. This contrasts with methods that fix a parametric form and use the reparameterization trick with Monte Carlo gradient estimates.
Finally, in Appendix I, we demonstrate that, by exploiting the Woodbury matrix identity, our inference updates require only $\mathcal{O}(m)$ time complexity ($m$ is the number of weights in the NN architecture). This linear scaling renders our Bayesian framework feasible for contemporary large-scale deep neural architectures, which are well suited to model high-dimensional data.
@@ 5 Experiments
Details on the experimental setup, including dataset descriptions, benchmark methods specifications and evaluation metrics definitions are provided in Appendix J. Moreover, the implementation details for NeuralSurv are provided in Appendix K.
To comprehensively evaluate NeuralSurv, we compare its performance against the following set of benchmark models: MTLR Yu2011, DeepHit Lee2018, DeepSurv Katzman2018, Logistic Hazard Gensheimer2019, CoxTime Kvamme2019, CoxCC Kvamme2019, PMF Kvamme2021, PCHazard Kvamme2021, BCESurv Kvamme2023, and DySurv Mesinovic2024, Sumo-Net rindt2022 and DQS Yanagisaw2023. A detailed overview of these models is provided in Appendix L and summarized in Table A2. Except for DySurv, which employs an autoencoder framework, we adopt the same NN architecture across all benchmark models and NeuralSurv to parameterize the hazard function. For DySurv, we use the original autoencoder architecture specified in its implementation.
We assess discriminative performance using the Antolini’s concordance index (C-index) Antolini2005, and evaluate model calibration with the inverse probability of censoring weighting (IPCW) integrated Brier score (IBS) Graf1999, the Distribution Calibration (D-Calibration) Haider2020, and the Kaplan–Meier Calibration (KM Calibration) Chapfuwa2023. The C-index evaluates how well a model performs by measuring the concordance between the rankings of the predicted event times and the true event times. The C-index ranges from 0 to 1, where higher values indicate better discriminative performance; a value of 0.5 corresponds to random guessing. Similar to the mean squared error, the Brier score (BS) assesses the accuracy of an estimated survival function at some time $t$. The IPCW are observation-specific weights that account for censoring in survival data, ensuring that the BS remains unbiased. The IPCW IBS is the integral of the IPCW BS over the observational period. The D‑Calibration test bins each individual’s predicted survival probability at their observed event time into equal-width bins over and applies a $\chi^{2}$ test to assess whether those predicted probabilities are uniform across bins. A well-calibrated model should yield a non-significant p‑value. The KM-Calibration procedure compares the average predicted survival curve with the Kaplan–Meier estimate. For the KM-Calibration, the closer the two curves align, the better calibrated the model, where 0 indicates perfect calibration and 1 indicates maximal miscalibration and a random prediction yields 0.25. The C-index and the IPCW IBS metrics are computed using the TorchSurv package Monod2024. The D-Calibration and KM-Calibration metrics are computed using the SurvivalEVAL package qi2024survivaleval.
@@ 5.1 Synthetic Data Experiment
In this section, we present experiments conducted on synthetic data. The experimental setup was inspired by teh_survival and constitutes a broadly applicable evaluation benchmark. We simulate the training sets with increasing sizes $N=25,50,100$ and $150$ samples where the event time was drawn from two distributions: $p_{0}(T)=\text{LogNormal}(3,0.8^{2})$ and $p_{1}(T)=\text{LogNormal}(3.5,1^{2})$. Each observation includes a covariate indicating whether the event time is sampled from $p_{0}$ or $p_{1}$, along with three additional noisy covariates generated from a standard normal distribution. The censoring times are drawn from an exponential distribution with a rate of $0.025$ yielding an average censoring rate of 54% across the four synthetic datasets. The test set is generated using the same data-generating process, fixed to $100$ observations, and held constant across all experiments.
Figure 1 presents the true survival function alongside the predicted functions from NeuralSurv and the two top-performing benchmark models, selected based on IPCW IBS. Each panel represents a different training set size. As the number of training samples increases, the predicted survival functions match more closely the true survival function. The results show that NeuralSurv consistently ranks as the best method according to IPCW IBS, and its predictive accuracy improves with larger sample sizes. Beyond its competitive performance, NeuralSurv also provides Bayesian credible intervals, offering uncertainty estimates for survival probabilities, an important feature absent in deep learning benchmark models. Notably, these credible intervals appropriately narrow as more data becomes available, demonstrating well-calibrated uncertainty quantification. Corresponding C-index, IPCW IBS score, D-Calibration p-values, and KM-Calibration scores for all methods are reported in Tables A3-A4.
@@ 5.2 Real Survival Data Experiments
Table 1 presents the C-index and IPCW IBS on the held-out test sets for three representative datasets, while results for the remaining datasets, as well as D-Calibration and KM-Calibration, are shown in Tables A5A6. Across eight datasets, NeuralSurv achieves the best IPCW IBS score on seven, highlighting superior overall calibration compared to benchmarks. It consistently passes D-calibration, together with Sumo-Net as the only benchmark achieving this result, while it ranks fifth in KM-calibration. This strength arises from its Bayesian framework, which naturally models uncertainty and provides effective regularization in data-scarce settings. Beyond calibration, NeuralSurv also demonstrates strong discriminative performance, achieving the best C-index in four datasets and the second best in three.
An ablation study using a larger training set of 250 observations is presented in Tables A7-A8. NeuralSurv continues to outperform benchmark methods under this setting in terms of calibration performance demonstrating the robustness of the method to training size. Furthermore, we also include results from traditional survival models, such as the Cox Proportional Hazards model Cox1972, the Weibull Accelerated Failure Time model Carroll2003, the Random Survival Forest Ishwaran2008, and the Survival Support Vector Machine Polsterl2015 in Tables A9-A10. These models often achieve strong performance in data-scarce regimes. However, they are not designed to leverage high-dimensional or complex feature representations, which limits their applicability in modern deep learning contexts. Our focus remains on evaluating deep survival methods that can scale with data complexity, but we include these classical baselines for reference and completeness. A prior sensitivity analysis for the parameter $\phi$, using priors with double and half the original variance, is presented in Table A11. While the posterior distributions under different priors largely overlapped, their central tendencies occasionally differed, indicating mild to moderate sensitivity to the choice of prior. Incorporating prior information remains important, as it helps strike a principled balance between model flexibility and regularization.
@@ 6 Conclusion
We propose the first fully Bayesian framework for deep survival analysis that models time-varying relationships between covariates and risk. On both synthetic and real-world datasets, in data-scarce regimes, our method consistently achieves better calibration than state-of-the-art deep survival models and matches or surpasses their discriminative performance. In contrast to previous approaches in deep survival analysis, which are either constrained to discrete-time settings Yu2011; Lee2018; Gensheimer2019; Kvamme2021; Kvamme2023; Mesinovic2024 or lack the ability to provide Bayesian uncertainty quantification Yu2011; Lee2018; Katzman2018; Gensheimer2019; Kvamme2019; Kvamme2023; Mesinovic2024, NeuralSurv introduces a continuous-time modeling framework that naturally incorporates Bayesian inference, enabling both accurate survival predictions and well-calibrated uncertainty estimates.
Pólya–Gamma and Poisson data-augmentation schemes (Section 3.3) have been extensively employed with standard Gaussian process models donner_opper_gaussian_cox; pg_poisson_hawkes. Likewise, the local linearization of Bayesian neural networks, which yields a Gaussian process–based approximation, (Section 4.2), is a well established technique Immer2021. To our knowledge, this work is the first to integrate these two approaches into a unified framework that capitalizes on Gaussian process conjugacy. By combining these methods, we contribute a novel inference strategy at the intersection of Bayesian deep learning and Gaussian process modeling.
Despite its strengths, NeuralSurv relies on three key simplifying assumptions. First, we assume a sigmoidal hazard function, a choice shared by prior work (e.g.,  teh_survival; Kim2018), which may not capture all risk dynamics. Second, our mean‑field variational inference treats parameters $\phi$ and $\boldsymbol{\theta}$ as independent, ignoring posterior correlations. Third, we linearize the network around the MAP estimate to enforce conjugacy. In real‑world settings, however, the true posterior can be multimodal and strongly correlated, so this local, factorized approximation may overlook secondary modes or misestimate joint uncertainty.
Concerning the computational efficiency of our method, the coordinate‑ascent updates scale linearly with network size but still require full‑dataset passes each iteration. For very large cohorts, this becomes a bottleneck. Extending the algorithm to use stochastic or mini‑batch updates would preserve conjugacy benefits while improving scalability.
@@
Acknowledgments and Disclosure of Funding
SB acknowledges funding from the MRC Centre for Global Infectious Disease Analysis (reference MR/X020258/1), funded by the UK Medical Research Council (MRC). This UK funded award is carried out in the frame of the Global Health EDCTP3 Joint Undertaking. SB acknowledges support from the Novo Nordisk Foundation via The Novo Nordisk Young Investigator Award (NNF20OC0059309). SB acknowledges the Danish National Research Foundation (DNRF160) through the chair grant. SB acknowledges support from The Eric and Wendy Schmidt Fund For Strategic Innovation via the Schmidt Polymath Award (G-22-63345) which also supports AM and MM.